OpenStudy (alyssastooks):

Evaluate the expression for the given value of the variable(s). -x 2 - 4x; x = –3 A. 3 B. –17 C. –1 D. 11 Simple math, medals given.

OpenStudy (tkhunny):

Rewrite your expression. Everywhere is says "x", substitute the value "-3". Simplify. Show your work.

OpenStudy (alyssastooks):

-3^2 - 4(-3);-3 = -3?

OpenStudy (tkhunny):

-3^2 - 4(-3) This is just fine, but one tiny point to discuss, It isn't always clear what to do with the minus sign out front when there is an exponent. Does -3^2 mean -(3^2) or does it mean (-3)^2. Always good to be perfectly clear. (-3)^2 - 4(-3) = 9 + 12 Ponder where those two values came from.

OpenStudy (alyssastooks):

I see what you mean but I haven't thought about any of this in 8 years, I'm not even sure how to simplify correctly. Do I simplify 9 + 12?

OpenStudy (alyssastooks):

I see how you came up with them, I remember most of the exponent rules but I'm helping my little sister in 6th grade.

OpenStudy (tkhunny):

No problem. Keep up the good work. Just keep simplifying. 9 + 12 = 21 That's all.

OpenStudy (alyssastooks):

So... 11...

OpenStudy (triciaal):